A class consists of 1/4 freshmen,1/5 sophomores, and 1/10
juniors; the rest are seniors. What fraction of the class is
seniors?

Answer by mananth(16946)

A class consists of 1/4 freshmen,1/5 sophomores, and 1/10
juniors; the rest are seniors. What fraction of the class is
seniors?
Add all the three
(1/4)+(1/5)+(1/10)
LCM = 20
(5+4+2)/20
=11/20
..
1-(11/20) will be seniors
LCM = 20
(20-11)/20
=9/20 are seniors
...
m.ananth@hotmail.ca